# SEDMamba: Enhancing Selective State Space Modelling with Bottleneck Mechanism and Fine-to-Coarse Temporal Fusion for Efficient Error Detection in Robot-Assisted Surgery

### v3

@@ Abstract
Automated detection of surgical errors can improve robotic-assisted surgery. Despite promising progress, existing methods still face challenges in capturing rich temporal context to establish long-term dependencies while maintaining computational efficiency. In this paper, we propose a novel hierarchical model named SEDMamba, which incorporates the selective state space model (SSM) into surgical error detection, facilitating efficient long sequence modelling with linear complexity. SEDMamba enhances selective SSM with a bottleneck mechanism and fine-to-coarse temporal fusion (FCTF) to detect and temporally localize surgical errors in long videos. The bottleneck mechanism compresses and restores features within their spatial dimension, thereby reducing computational complexity. FCTF utilizes multiple dilated 1D convolutional layers to merge temporal information across diverse scale ranges, accommodating errors of varying duration. Our work also contributes the first-of-its-kind, frame-level, in-vivo surgical error dataset to support error detection in real surgical cases. Specifically, we deploy the clinically validated observational clinical human reliability assessment tool (OCHRA) to annotate the errors during suturing tasks in an open-source radical prostatectomy dataset (SAR-RARP50). Experimental results demonstrate that our SEDMamba outperforms state-of-the-art methods with at least 1.82% AUC and 3.80% AP performance gains with significantly reduced computational complexity. The corresponding error annotations, code and models will be released at https://github.com/wzjialang/SEDMamba.
@@
Surgical error detection, Robotic-assisted surgery, Selective state space model, Video, Temporal
@@ I Introduction
  1. 1.
    We propose a novel hierarchical model, SEDMamba, constructed with bottleneck multi-scale state space (BMSS) blocks to detect surgical errors in long videos from real RAS cases. To handle errors with variable duration, we design a fine-to-coarse temporal fusion (FCTF) using different dilated 1D convolutional layers to merge temporal information across different scale ranges. Besides, to further reduce computational complexity, we incorporate a bottleneck mechanism that compresses the spatial dimension of features, enabling selective SSM to concentrate on modelling the temporal dimension. To the best of our knowledge, this is the first attempt to exploit selective SSM in SED.
  2. 2.
    We present the first-of-its-kind, frame-level surgical error annotations of an in-vivo RAS dataset and apply them to train and validate the proposed SEDMamba model. Our work establishes a roadmap of how to utilize this new dataset for developing SED benchmarks.
  3. 3.
    The proposed SEDMamba outperforms state-of-the-art methods by 1.82-13.06% AUC and 3.8-18.86% AP while requiring 11.71-91.67% fewer model parameters and 40.71-92.32% fewer FLOPs.
II Methodology
II-A Dataset Description and Annotation
@@ II-B Surgical Error Detection Model (SEDMamba)
Figure 3: The overall architecture of the proposed SEDMamba. (a) The fundamental block of SEDMamba, namely the bottleneck multi-scale state space (BMSS) block; (b) Fine-to-coarse temporal fusion (FCTF); (c) Bottleneck mechanism.
@@ II-B3 Bottleneck Multi-scale State Space (BMSS) Block
The structure of the BMSS block is shown in Fig. 3(a). Encouraged by the strong long-range modelling ability of SSMs, the BMSS block incorporates selective SSM with bottleneck mechanism and fine-to-coarse temporal fusion (FCTF) to reduce computational complexity and capture inter-frame long-range dependencies across multiple temporal scales based on spatial embedding sequences, respectively. Given input $x\in\mathbb{R}^{L\times D}$, it first undergoes a bottleneck 1D convolution to compress its spatial dimension from $D$ to $G$, where compression factor $G\ll D$. Multi-scale temporal information is then obtained and merged via FCTF across four different temporal scales to generate fine-to-coarse temporal feature $f_{FC}$. After that, the $f_{FC}$ is expanded with a linear layer and split into two information flows. The first flow passes through a 1D convolutional layer, followed by a SiLU activation function [34] before entering the selective SSM. The output of the selective SSM is then gated by the $z$ generated from the other flow. Finally, we use another bottleneck 1D convolutional layer to restore the spatial dimension to $\frac{D}{2}$ and get the output of the BMSS block $y\in\mathbb{R}^{L\times\frac{D}{2}}$. Formally, we present the operations of the BMSS block in Algorithm 1. The BMSS block inherits the linear complexity of the selective SSM while retaining a local-to-global receptive field, which is promising for video-based temporal error detection.
@@ II-B4 Fine-to-Coarse Temporal Fusion (FCTF)
Algorithm 1 Bottleneck Multi-scale State Space (BMSS) Block
[…]
2:Output: $y:(L,D/2)$.
3:/* Bottleneck: Compress $x$ along spatial dimension */
4:$f_{C}:(L,G)\leftarrow\text{1DConv}^{G}(x)$
5:/* Fine-to-Coarse Temporal Fusion (FCTF): Obtain
6:    multi-scale temporal information across four scales*/
7:$f_{T1}:(L,E)\leftarrow\text{D1DConv}^{E}(f_{C})$ with 2 dilation rate
8:$f_{T2}:(L,E)\leftarrow\text{D1DConv}^{E}(f_{T}1)$ with 4 dilation rate
9:$f_{T3}:(L,E)\leftarrow\text{D1DConv}^{E}(f_{T}2)$ with 8 dilation rate
10:$f_{FC}:(L,G)\leftarrow\text{1DConv}^{G}\left(\text{Cat}(f_{C},f_{T1},f_{T2},f_{T3})\right)$
11:/* Information split */
12:$x:(L,2G)\leftarrow\text{Linear}^{2G}(f_{FC})$
13:$x,z:(L,G)\leftarrow\text{Split}(x)$
[…]
15:/* Selective SSM*/
16:$\mathbf{A}:(G,N)\leftarrow\text{Parameter}(\mathbf{A})$
17:$\mathbf{B}:(L,N)\leftarrow\text{Linear}_{\mathbf{B}}^{N}(x)$
18:$\mathbf{C}:(L,N)\leftarrow\text{Linear}_{\mathbf{C}}^{N}(x)$
[…]
20:                  $\left.\left.\text{Broadcast}^{G}(\text{Linear}_{\Delta}(x))\right)\right)$
21:$\overline{\mathbf{A}},\overline{\mathbf{B}}:(L,G,N)\leftarrow\text{Discretization}(\Delta,\overline{\mathbf{A}},\overline{\mathbf{B}})$
22:$x:(L,G)\leftarrow\text{SSM}(\overline{\mathbf{A}},\overline{\mathbf{B}},\mathbf{C})(x)$
23:/* Get gated $y$/
24:$y:(L,G)\leftarrow\text{Linear}^{G}(x\otimes\text{SiLU}(z))$
25:/* Bottleneck: Restore $y$ along spatial dimension */
26:$y:(L,D/2)\leftarrow\text{1DConv}^{D/2}(y)$
27:return $y:(L,D/2)$
@@ III-A Experimental Settings
All experiments are implemented in PyTorch on a RTX 4090 GPU. We use AdamW optimizer with initial learning rate of 1e-4 and maximum training epoch set at 100. We sampled SAR-RARP videos at 5Hz to get 40 videos with 46,586 normal frames and 18,540 error frames as the training set, and 8 videos with 7549 normal frames and 2096 error frames as the testing set. We use DINOv2 with fixed weights pretrained on ImageNet [33] as the visual extractor to generate spatial embedding sequences $\mathcal{E}$ from the video with $L$ frames as the model input. The compression factor $G$ in the BMSS block is set as 64. The channel $E$ of dilated 1D convolutional layers in FCTF is fixed as $\frac{D}{8}$. Following  [27], the state dimension $N$ of selective SSM is set to $16$.
@@ III-B Evaluation Metrics
TABLE II: Quantitative results of SEDMamba and other methods for all test data and different error duration ranges. Note that all metrics are calculated on frame level.
AUC (%) $\uparrow$ AP (%) $\uparrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
LSTM [22] 68.43$\pm$1.21 37.26$\pm$2.65 74.36$\pm$0.90 33.06$\pm$1.27 61.54$\pm$0.62 21.34$\pm$1.56
TeCNO [35] 58.14$\pm$0.33 26.01$\pm$0.32 57.95$\pm$1.90 22.23$\pm$1.41 58.72$\pm$0.90 16.00$\pm$0.20
MS-TCN [36] 58.82$\pm$0.16 29.98$\pm$3.06 55.73$\pm$1.29 29.06$\pm$0.39 61.26$\pm$2.94 14.93 $\pm$1.10
MS-TCN++ [37] 58.57$\pm$0.43 30.35$\pm$1.35 55.34$\pm$1.26 26.48$\pm$1.95 61.44$\pm$0.48 17.99$\pm$1.53
ASFormer [24] 68.76$\pm$0.93 36.79$\pm$2.35 75.07$\pm$1.34 32.70$\pm$2.24 61.72$\pm$2.24 20.14$\pm$0.66
Mamba [27] 69.38$\pm$0.26 41.07$\pm$0.95 75.73$\pm$0.42 35.12$\pm$1.72 62.41$\pm$0.74 25.10$\pm$1.78
Vim [28] 69.06$\pm$1.46 40.63$\pm$2.30 75.24$\pm$0.88 34.64$\pm$1.54 62.02$\pm$2.30 24.67$\pm$1.08
SEDMamba (Ours) 71.20$\pm$0.26 44.87$\pm$1.52 77.92$\pm$0.76 38.91$\pm$0.57 64.01$\pm$0.85 26.71$\pm$0.50
@@
TABLE IV: Ablation results for bottleneck mechanism and FCFT in BMSS blocks.
Bottleneck FCFT Params (K) $\downarrow$ FLOPs (M) $\downarrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
@@ G Params (K) $\downarrow$ FLOPs (M) $\downarrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
III-C Comparison with the State-of-the-Art
We compare the proposed SEDMamba with state-of-the-art SED method: LSTM [22]; convolutional and Transformer-based video analysis methods: TeCNO [35], MS-TCN [36], MS-TCN++ [37], ASFormer [24]; and selective SSM-based methods: Mamba [27] and Vim [28]. All methods are implemented based on their released code and original literature, and fine-tuned to fit our error detection task to provide a basis for a fair comparison.
Table II shows quantitative results of all methods. For the overall performance (i.e. all test data regardless of error duration), the proposed SEDMamba consistently achieves superior performance with an increase of 1.82-13.06% in AUC and 3.8-18.86% in AP, compared to other methods. Notably, all selective SSM-based methods surpass the Transformer-based method ASFormer, demonstrating the superiority of selective SSM in the SED task. Moreover, SEDMamba achieves gains of 2.19-22.58% in AUC and 3.79-16.68% in AP for errors with short duration ($<$ 3 seconds), and 1.6-5.29% in AUC and 1.61-11.78% in AP for errors with long duration ($>$ 3 seconds).
Fig. 4 illustrates typical error detection results on test videos. [^1] Additional detection results via various error duration are included in the supplementary video. The second-best method Mamba tends to output smoother and more ambiguous prediction curves, which impedes its sensitivity to errors. It also produces lower confidence outputs within error frames. Thanks to the fine-to-coarse temporal information provided by FCTF for selective SSM, SEDMamba is able to identify suddenly occurring errors with short duration (highlighted in red boxes) and consistently produces high confidence within error areas in most cases (highlighted in black boxes). In most normal frames, SEDMamba has comparable or lower error probabilities to Mamba, with only a small portion showing higher error probabilities. Note that both AUC and AP comprehensively consider all classes, i.e. ‘error’ and ‘normal’, indicating SEDMamba has superior performance in both classes.
III-D Computational Efficiency
@@ III-E1 Effectiveness of Key Components in BMSS Block
We study the effect of bottleneck mechanism and FCFT in BMSS blocks. As shown in Table IV, whereby removing each component leads to performance degradation. The bottleneck mechanism yields performance gains of +1.64% in AUC and +3.69% in AP, while model parameters and FLOPs are significantly decreased by 82.98% and 78.83%, which guards against overfitting and enhances model efficiency. Besides, such a bottleneck mechanism compresses features on the spatial dimension, which can mitigate the impact of redundant spatial noise on long-term dependency modelling. Introducing FCTF to bottleneck mechanism leads to a slight increase in computational complexity, which is acceptable given the substantial performance gains of +1.58% in AUC and +3.58% in AP. This underscores the importance of providing multi-scale temporal information for selective SSM in the SED task.

### v4

@@ Abstract
Automated detection of surgical errors can improve robotic-assisted surgery. Despite promising progress, existing methods still face challenges in capturing rich temporal context to establish long-term dependencies while maintaining computational efficiency. In this paper, we propose a novel hierarchical model named SEDMamba, which incorporates the selective state space model (SSM) into surgical error detection, facilitating efficient long sequence modelling with linear complexity. SEDMamba enhances selective SSM with a bottleneck mechanism and fine-to-coarse temporal fusion (FCTF) to detect and temporally localize surgical errors in long videos. The bottleneck mechanism compresses and restores features within their spatial dimension, thereby reducing computational complexity. FCTF utilizes multiple dilated 1D convolutional layers to merge temporal information across diverse scale ranges, accommodating errors of varying duration. Our work also contributes the first-of-its-kind, frame-level, in-vivo surgical error dataset to support error detection in real surgical cases. Specifically, we deploy the clinically validated observational clinical human reliability assessment tool (OCHRA) to annotate the errors during suturing tasks in an open-source radical prostatectomy dataset (SAR-RARP50). Experimental results demonstrate that our SEDMamba outperforms state-of-the-art methods with at least 1.82% AUC and 3.80% AP performance gains with significantly reduced computational complexity. The corresponding error annotations, code and models are released at https://github.com/wzjialang/SEDMamba.
@@
Computer vision for medical robotics, surgical robotics: laparoscopy, data sets for robotic vision, surgical error detection, selective state space model
publicationid: pubid:  © 2024 IEEE. Personal use is permitted, but republication/redistribution requires IEEE permission. See https://www.ieee.org/publications/rights/index.html for more information
@@ I Introduction
1) We propose a novel hierarchical model, SEDMamba, constructed with bottleneck multi-scale state space (BMSS) blocks to detect surgical errors in long videos from real RAS cases. To handle errors with variable duration, we design a fine-to-coarse temporal fusion (FCTF) using different dilated 1D convolutional layers to merge temporal information across different scale ranges. Besides, to further reduce computational complexity, we incorporate a bottleneck mechanism that compresses the spatial dimension of features, enabling selective SSM to concentrate on modelling the temporal dimension. To the best of our knowledge, this is the first attempt to exploit selective SSM in SED.
2) We present the first-of-its-kind, frame-level surgical error annotations of an in-vivo RAS dataset and apply them to train and validate the proposed SEDMamba model. Our work establishes a roadmap of how to utilize this new dataset for developing SED benchmarks.
3) The proposed SEDMamba outperforms state-of-the-art methods by 1.82-13.06% AUC and 3.8-18.86% AP while requiring 11.71-91.67% fewer model parameters and 40.71-92.32% fewer FLOPs.
@@ II-B Surgical Error Detection Model (SEDMamba)
Figure 3: The pipeline of the proposed SEDMamba. (a) The overall architecture of SEDMamba. (b) The fundamental block of SEDMamba, namely the bottleneck multi-scale state space (BMSS) block; (c) Fine-to-coarse temporal fusion (FCTF); (d) Bottleneck mechanism.
@@ II-B3 Bottleneck Multi-scale State Space (BMSS) Block
The structure of the BMSS block is shown in Fig. 3(a). Encouraged by the strong long-range modelling ability of SSMs, the BMSS block incorporates selective SSM with bottleneck mechanism and fine-to-coarse temporal fusion (FCTF) to reduce computational complexity and capture inter-frame long-range dependencies across multiple temporal scales based on spatial embedding sequences, respectively. Given input $x\in\mathbb{R}^{L\times D}$, it first undergoes a bottleneck 1D convolution to compress its spatial dimension from $D$ to $G$, where compression factor $G\ll D$. Multi-scale temporal information is then obtained and merged via FCTF across four different temporal scales to generate fine-to-coarse temporal feature $f_{FC}$. After that, the $f_{FC}$ is expanded with a linear layer and split into two information flows. The first flow passes through a 1D convolutional layer, followed by a SiLU activation function [34] before entering the selective SSM. The output of the selective SSM is then gated by the $z$ generated from the other flow. Finally, we use another bottleneck 1D convolutional layer to restore the spatial dimension to $\frac{D}{2}$ and get the output of the BMSS block $y\in\mathbb{R}^{L\times\frac{D}{2}}$.  [^1] Pseudo-code of the BMSS block is provided as Algorithm 1 in the supplementary video. The BMSS block inherits the linear complexity of the selective SSM while retaining a local-to-global receptive field, which is promising for video-based temporal error detection.
@@ III-A Experimental Settings
All experiments are implemented in PyTorch on a RTX 4090 GPU. We use AdamW optimizer with initial learning rate of 1e-4 and maximum training epoch set at 100. We sampled SAR-RARP videos at 5Hz to get 40 videos with 46,586 normal frames and 18,540 error frames as the training set, and 8 videos with 7549 normal frames and 2096 error frames as the testing set. We use DINOv2 (giant version with registers) with fixed weights pretrained on ImageNet [33] as the visual extractor to generate spatial embedding sequences $\mathcal{E}$ from the video with $L$ frames as the model input. The compression factor $G$ in the BMSS block is set as 64. The channel $E$ of dilated 1D convolutional layers in FCTF is fixed as $\frac{D}{8}$. Following  [27], the state dimension $N$ of selective SSM is set to $16$.
@@ III-B Evaluation Metrics
III-C Comparison with the State-of-the-Art
We compare the proposed SEDMamba with state-of-the-art SED method: LSTM [22]; convolutional and Transformer-based video analysis methods: TeCNO [35], MS-TCN [36], MS-TCN++ [37], ASFormer [24]; and selective SSM-based methods: Mamba [27] and Vim [28]. All methods are implemented based on their released code and original literature, and fine-tuned to fit our error detection task to provide a basis for a fair comparison. Note that we use an original Mamba/Vim block and a 1D convolutional layer with 1 kernel size and 1 channel to adapt Mamba [27]/Vim [28] to the SED task.
Table II shows quantitative results of all methods. For the overall performance (i.e. all test data regardless of error duration), the proposed SEDMamba consistently achieves superior performance with an increase of 1.82-13.06% in AUC and 3.8-18.86% in AP, compared to other methods. Notably, all selective SSM-based methods surpass the Transformer-based method ASFormer, demonstrating the superiority of selective SSM in the SED task. SEDMamba also outperforms other methods when grouping annotations together and detecting error instances instead of individual frames. Moreover, SEDMamba achieves gains of 2.19-18.68% in AUC and 3.79-21.76% in AP for short errors ($<$ 3 seconds), and 1.6-11.02% in AUC and 1.61-13.78% in AP for long errors ($\geq$ 3 seconds).
Fig. 4 illustrates typical error detection results on test videos. [^2] Additional detection results via various error duration are included in the supplementary video. The second-best method Mamba tends to output smoother and more ambiguous prediction curves, which impedes its sensitivity to errors. It also produces lower confidence outputs within error frames. Thanks to the fine-to-coarse temporal information provided by FCTF for selective SSM, SEDMamba is able to identify suddenly occurring errors with short duration (highlighted in red boxes) and consistently produces high confidence within error areas in most cases (highlighted in black boxes). In most normal frames, SEDMamba has comparable or lower error probabilities to Mamba, with only a small portion showing higher error probabilities. Specifically, when video quality deteriorates, such as instruments being covered with blood (second example in the 2nd row of Fig. 4) or affected by glare (second example in the 3rd row of Fig. 4), both SEDMamba and Mamba may produce false positives, with SEDMamba tending to output higher error probabilities. Note that both AUC and AP comprehensively consider all classes, i.e. ‘error’ and ‘normal’, indicating SEDMamba has superior performance in both classes.
TABLE II: Quantitative results of SEDMamba and other methods for all test data and different error duration ranges. We calculate AUC and AP at individual frame-level and per instance of error (AUCins and APins). Instance-level calculations are done by grouping consecutive frames with the same label as an error instance and then averaging the predicted error probabilities, across all frames within an instance, to obtain the overall error probability prediction per instance.
AUC (%) $\uparrow$ AP (%) $\uparrow$ AUCins (%) $\uparrow$ APins (%) $\uparrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
LSTM [22] 68.43$\pm$1.21 37.26$\pm$2.65 56.58$\pm$0.83 60.35$\pm$0.93 74.36$\pm$0.90 33.06$\pm$1.27 61.54$\pm$0.62 21.34$\pm$1.56
TeCNO [35] 58.14$\pm$0.33 26.01$\pm$0.32 51.87$\pm$0.56 52.06$\pm$0.61 59.24$\pm$0.36 17.15$\pm$0.44 57.65$\pm$0.85 14.06$\pm$0.37
MS-TCN [36] 58.82$\pm$0.16 29.98$\pm$3.06 51.17$\pm$0.86 50.36$\pm$1.69 62.50$\pm$1.23 22.76$\pm$5.69 54.70$\pm$1.68 14.37$\pm$0.46
MS-TCN++ [37] 58.57$\pm$0.43 30.35$\pm$1.35 51.69$\pm$0.12 51.22$\pm$0.19 63.57$\pm$0.87 24.75$\pm$1.85 52.99$\pm$0.06 12.93$\pm$0.46
ASFormer [24] 68.76$\pm$0.93 36.79$\pm$2.35 54.92$\pm$0.87 56.38$\pm$2.17 74.77$\pm$2.02 25.59$\pm$9.12 62.66$\pm$2.74 17.99$\pm$2.46
Mamba [27] 69.38$\pm$0.26 41.07$\pm$0.95 57.21$\pm$2.14 62.24$\pm$1.45 75.73$\pm$0.42 35.12$\pm$1.72 62.41$\pm$0.74 25.10$\pm$1.78
Vim [28] 69.06$\pm$1.46 40.63$\pm$2.30 57.76$\pm$1.25 62.02$\pm$2.53 75.24$\pm$0.88 34.64$\pm$1.54 62.02$\pm$2.30 24.67$\pm$1.08
SEDMamba (Ours) 71.20$\pm$0.26 44.87$\pm$1.52 59.28$\pm$0.36 63.95$\pm$0.57 77.92$\pm$0.76 38.91$\pm$0.57 64.01$\pm$0.85 26.71$\pm$0.50
@@ III-C Comparison with the State-of-the-Art
III-D Computational Efficiency
@@
TABLE IV: Ablation results for bottleneck mechanism and FCTF in BMSS blocks.
Bottleneck FCTF Params (K) $\downarrow$ FLOPs (M) $\downarrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
@@ G Params (K) $\downarrow$ FLOPs (M) $\downarrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
TABLE VII: Comparison of different numbers of dilated convolution layers in FCTF.
Numbers [Dilation Rates] Params (K) $\downarrow$ FLOPs (M) $\downarrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
0 267.11 22.66 69.62$\pm$0.65 41.29$\pm$0.84
1 [2] 285.76 26.35 69.85$\pm$0.74 41.37$\pm$0.44
2 [2, 4] 287.90 26.57 69.89$\pm$0.56 41.77$\pm$0.80
3 [2, 4, 8] 290.03 26.78 71.20$\pm$0.26 44.87$\pm$1.52
4 [2, 4, 8, 16] 292.17 26.99 70.45$\pm$0.90 43.36$\pm$2.50
5 [2, 4, 8, 16, 32] 294.30 27.20 70.30$\pm$1.57 42.33$\pm$1.98
TABLE VIII: Comparison of Mamba alternatives in BMSS blocks.
Methods Params (K) $\downarrow$ FLOPs (M) $\downarrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
BMSSLSTM 291.95 29.22 68.78$\pm$0.06 41.58$\pm$1.65
BMSSTeCNO 600.88 19.08 67.19$\pm$0.26 39.84$\pm$0.23
BMSSASFormer 354.55 35.28 66.83$\pm$1.03 34.70$\pm$1.88
BMSS (our SEDMamba) 290.03 26.78 71.20$\pm$0.26 44.87$\pm$1.52
@@ III-E1 Effectiveness of Key Components in BMSS Block
We study the effect of bottleneck mechanism and FCTF in BMSS blocks. As shown in Table IV, whereby removing each component leads to performance degradation. The bottleneck mechanism yields performance gains of +1.64% in AUC and +3.69% in AP, while model parameters and FLOPs are significantly decreased by 82.98% and 78.83%, which guards against overfitting and enhances model efficiency. Besides, such a bottleneck mechanism compresses features on the spatial dimension, which can mitigate the impact of redundant spatial noise on long-term dependency modelling. Introducing FCTF to bottleneck mechanism leads to a slight increase in computational complexity, which is acceptable given the substantial performance gains of +1.58% in AUC and +3.58% in AP. This underscores the importance of providing multi-scale temporal information for selective SSM in the SED task.